OpenStudy (anonymous):

what are the coordinates of the point on the line 7x-3y=11 that is closest to the origin? should i just let x= 0 and solve for y and then let y=0 and solve for x and get (x,y)? a) (1.57, 0) b) (1.33, 0.47) c) (1.33. -0.57) d) (1.43, -0.47) e) (1.27, -0.67)

OpenStudy (dumbcow):

find the equation of line perpendicular to given line that goes through origin. point where these 2 lines intersect is answer

myininaya (myininaya):

d=sqrt((x-0)^2+(y-0)^2) where (x,y) is a point on y=7x/3-11/3 d=sqrt(x^2+y^2) d=sqrt(x^2+(7x/3-11/3)^2) d^2=x^2+(7x/3-11/3)^2 f(x)=x^2+(7x/3-11/3)^2 find f'(x) to minimize

myininaya (myininaya):

if x=77/58, then y=(7/3)*(77/58)-11/3